# Distributional Finite Element curl div Complexes and Application to Quad Curl Problems

### v3

@@ Abstract.
The paper addresses the challenge of constructing conforming finite element spaces for high-order differential operators in high dimensions, with a focus on the ${\rm curl\,}\operatorname{div}$ operator in three dimensions. Tangential-normal continuity is introduced in order to develop distributional finite element ${\rm curl\,}\operatorname{div}$ complexes. The spaces constructed are applied to discretize a quad curl problem, demonstrating optimal order of convergence. Furthermore, a hybridization technique is proposed, demonstrating its equivalence to nonconforming finite elements and weak Galerkin methods.
@@ 1. Introduction
In this work, we will design tangential-normal continuous finite element spaces for the $2$nd order differential operator curldiv and construct the corresponding distributional finite element curldiv complexes in three dimensions.
@@
For the discretization of the biharmonic equation in two dimensions, the so-called Hellan-Herrmann-Johnson (HHJ) mixed method [22, 23, 31] requires only normal-normal continuous finite elements for symmetric tensors and thus there is no need to use $C^{1}$-conforming elements for displacement. This normal-normal continuous finite element is then employed to solve linear elasticity [38] and Reissner-Mindlin plates [39], and used to construct the first two-dimensional distributional finite element divdiv complexes in [11]. Recently, the distributional finite element divdiv element for solving the mixed formulation of the biharmonic equation has been extended to arbitrary dimensions in [14].
Now we move to the ${\rm curl\,}\operatorname{div}$ operator between Sobolev spaces of negative order
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\xrightarrow{{\rm curl\,}\operatorname{div}}H^{-1}(\operatorname{div},\Omega),$
where $\mathbb{T}$ is the space of traceless tensors, $H^{-1}(\Omega):=(H_{0}^{1}(\Omega))^{\prime}$, and
We will construct tangential-normal continuous finite element spaces and apply to solve the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$.
Finite element spaces conforming to
$H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T}):=\left\{\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{T}):{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})\right\}$
are relatively complicated due to the smoothness requirement ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})$. While in the distributional sense
@@
Now the function $\boldsymbol{\tau}$ is tangential-normally continuous and $\boldsymbol{v}$ is only tangentially continuous so that $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=\operatorname*{rot}_{F}\boldsymbol{v}$ is continuous on face $F$. One can easily show $({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}$ in the distribution sense by taking $\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3})$ in (1).
@@
where $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ denote two mutually perpendicular unit tangential vectors of face $F$ and are used to determine the tangential component of the vector $\boldsymbol{\tau}\boldsymbol{n}$. The global finite element space $\Sigma_{k,h}^{\rm tn}$ by requiring single valued (2a) is tangential-normally continuous. A different tangential-normally continuous finite element in [21, 33] requires $(\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})|_{F}\in\mathbb{P}_{k-1}(F)$. That is the tangential-normal trace is one degree lower. A bubble-enriched space of $\Sigma_{k,h}^{\rm tn}$ is designed in [20] for Stokes equation with weakly imposed stress symmetry.
We use Nédélec elements $\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\subset V_{0}^{{\rm curl\,}}$ for the tangential continuous vector space. Use the Riesz representation of the $L^{2}$-inner product to bring the abstract dual to a concrete function. Define $({\rm curl\,}\operatorname{div})_{h}:\Sigma_{h}^{\rm tn}\to\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ such that
$(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h}\rangle,\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},$
@@
$\mathbb{R}^{3}\times\{0\}\rightarrow\mathbb{V}_{k+1,h}^{{\rm grad\,}}(\mathbb{R}^{3})\times\mathbb{R}$ $\xrightarrow{({\rm grad\,}\!,\,\operatorname{mskw}\boldsymbol{x})}\mathbb{V}_{k,h}^{{\rm curl\,}}(\mathbb{M})\xrightarrow{\operatorname{dev}{\rm curl\,}}$
@@
where $({\rm grad\,},\operatorname{mskw}\boldsymbol{x})\begin{pmatrix}\boldsymbol{v}\\ c\end{pmatrix}={\rm grad\,}\boldsymbol{v}+c\operatorname{mskw}\boldsymbol{x}$, and $\ell=k-1$ or $\ell=k$ is introduced to distinguish the first and second kind of Nédélec element. It is a discretization of the distributional ${\rm curl\,}\operatorname{div}$ complex:
@@
The lowest order, i.e. $k=1,\ell=0$, of the last three elements are illustrated in Fig. 1.
@@
(3) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{k-1,h},\psi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}},$
(4) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}},$
where $b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v}):=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v}).$
We prove two discrete inf-sup conditions and thus obtain the well-posedness of (3)-(4) and optimal order convergence
@@
Furthermore, we apply hybridization techniques to (3)-(4) which is equivalent to a stabilization-free weak Galerkin method and to the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements in [29, 47] for solving the quad curl problem. For other discretization of the quad-curl problem, we refer to the macro finite element method in [28], nonconforming finite element methods in [30, 41, 43], mixed finite element methods in [40, 10], decoupled finite element methods in [8, 46, 7], and references cited therein.
@@ 2. Distributional ${\rm curl\,}\operatorname{div}$ complex
In this section, we present the distributional ${\rm curl\,}\operatorname{div}$ complex. In particularly, we introduce the weak differential operator $({\rm curl\,}\operatorname{div})_{w}$ which can be defined on the tangential-normal continuous matrix functions.
@@ 2.1. Notation
Denote the space of all $3\times 3$ matrices by $\mathbb{M}$, and all trace-free/traceless $3\times 3$ matrices by $\mathbb{T}$. Define $\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\frac{1}{3}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I}$. Obviously for a scalar function $u$, $\operatorname{dev}(u\boldsymbol{I})=0$. For a vector $\boldsymbol{w}=(\omega_{1},\omega_{2},\omega_{3})^{\intercal}\in\mathbb{R}^{3}$, let
@@ 2.2. The ${\rm curl\,}\operatorname{div}$ complexes
Next we follow the framework developed in [12] to present a Helmholtz decomposition of $H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$. Denote by
@@
Notice that only interior faces are included in the second term as $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=0$ for $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega)$.
@@ 3.1. Finite element spaces
$\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T})=\mathbb{P}_{k-1}(T)\otimes\textrm{span}\{\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{i}\otimes\boldsymbol{t}_{i\ell}),\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{j}\otimes\boldsymbol{t}_{j\ell}),\ell=0,\ldots,3\}.$
@@
The construction can be readily extended to arbitrary dimension $\mathbb{R}^{d}$. There are $d+1$ faces for a $d$-simplex. At each face, we have $d-1$ linearly independent traceless matrices $\{\boldsymbol{t}_{F,i}\otimes\boldsymbol{n}_{F}\}_{i=1}^{d-1}$ and in total $(d+1)(d-1)$ such matrices form a basis of $\mathbb{T}$. The basis of bubble functions is constructed vertices-wisely $\{\lambda_{\ell}\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell})\}$ for $\ell=0,\ldots,d$ and $d-1$ different $i$ for each $\ell$. ∎
@@ 3.2. Distributional finite element complex
Notice that $\operatorname{div}_{h}$ is the $L^{2}$-adjoint of $-{\rm grad\,}$ restricted to $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Again for a fixed triangulation $\mathcal{T}_{h}$, $\operatorname{div}_{h}$ is well defined which can be obtained by inverting the mass matrix of $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. However $\{\operatorname{div}_{h}\}$ is not uniformly bounded as $h\to 0$ as $(\boldsymbol{v},{\rm grad\,}\psi)$ is in general not a bounded linear functional of $L^{2}$ unless $\boldsymbol{v}\in H^{1}(\Omega)$.

### v4

@@ Abstract.
The paper addresses the challenge of constructing conforming finite element spaces for the ${\rm curl\,}\operatorname{div}$ operator in three dimensions. Tangential-normal continuity is introduced in order to develop distributional finite element ${\rm curl\,}\operatorname{div}$ complexes. The spaces constructed are applied to discretize the quad curl problem, demonstrating optimal order of convergence. Furthermore, a hybridization technique is proposed, demonstrating its equivalence to nonconforming finite elements and weak Galerkin methods.
@@ 1. Introduction
In this work, we will design tangential-normal continuous finite element spaces for the $2$nd order differential operator curldiv and construct the corresponding distributional finite element curldiv complexes in three dimensions. We will apply the constructed finite elements to solve the fourth-order curl problem $-{\rm curl\,}\Delta{\rm curl\,}\boldsymbol{u}=\boldsymbol{f},\operatorname{div}\boldsymbol{u}=0$ in a domain $\Omega\subset\mathbb{R}^{3}$ with boundary conditions $\boldsymbol{u}\times\boldsymbol{n}=\operatorname{curl}\boldsymbol{u}=0$ on $\partial\Omega$. Such problem arises from multiphysics simulation such as modeling a magnetized plasma in magnetohydrodynamics [5, 33, 10].
@@
For the discretization of the biharmonic equation in two dimensions, the so-called Hellan-Herrmann-Johnson (HHJ) mixed method [23, 24, 32] requires only normal-normal continuous finite elements for symmetric tensors and thus $C^{0}$-conforming Lagrange element, not $C^{1}$-conforming elements, can be used for displacement. This normal-normal continuous finite element is then employed to solve linear elasticity [39] and Reissner-Mindlin plates [40], and used to construct the first two-dimensional distributional finite element divdiv complexes in [12]. Recently, the distributional finite element divdiv element for solving the mixed formulation of the biharmonic equation has been extended to arbitrary dimensions in [17].
Now we move to the curl div operator. Introduce the space $H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T}):=\left\{\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{T}):{\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})\right\}$, where $\mathbb{T}$ is the space of traceless tensors. A mixed formulation of the quad-curl problem is to find $\boldsymbol{\sigma}\in H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$, $\boldsymbol{u}\in L^{2}(\Omega;\mathbb{R}^{3})$ and $\phi\in H_{0}^{1}(\Omega)$ such that
$(\boldsymbol{\sigma},\boldsymbol{\tau})+b(\boldsymbol{\tau},\psi;\boldsymbol{u})$ $=0,$ $\boldsymbol{\tau}\in H(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega),$
$b(\boldsymbol{\sigma},\phi;\boldsymbol{v})$ $=-\langle\boldsymbol{f},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in L^{2}(\Omega;\mathbb{R}^{3}),$
where the bilinear form $b(\boldsymbol{\tau},\psi;\boldsymbol{v}):=(\operatorname{curl}{\operatorname{div}}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v}).$ The term $({\rm grad\,}\psi,\boldsymbol{u})$ is introduced to impose the divergence free condition $\operatorname{div}\boldsymbol{u}=0$.
Finite element spaces conforming to $H({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$ are relatively complicated due to the smoothness requirement ${\rm curl\,}\operatorname{div}\boldsymbol{\tau}\in L^{2}(\Omega;\mathbb{R}^{3})$. In the distributional sense
@@
(1) $\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau},\boldsymbol{v}\rangle:=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},{\rm curl\,}\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}(\llbracket\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\rrbracket,\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v})_{F},$
which is analog to the weak divdiv operator in HHJ mixed method [23, 24, 32]. Now the function $\boldsymbol{\tau}$ is tangential-normally continuous and $\boldsymbol{v}$ is tangentially continuous so that $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=\operatorname*{rot}_{F}\boldsymbol{v}$ is continuous on face $F$. One can easily show $({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}={\rm curl\,}\operatorname{div}\boldsymbol{\tau}$ in the distribution sense by taking $\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{3})$ in (1).
@@
where $\boldsymbol{t}_{1}$ and $\boldsymbol{t}_{2}$ denote two mutually perpendicular unit tangential vectors of face $F$ and are used to determine the tangential component of the vector $\boldsymbol{\tau}\boldsymbol{n}$. The global finite element space $\Sigma_{k,h}^{\rm tn}$ by requiring single valued (2a) is tangential-normally continuous. A different tangential-normally continuous finite element proposed in [22, 34] requires that $(\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})|_{F}\in\mathbb{P}_{k-1}(F)$, meaning the tangential-normal trace is reduced by one degree. A bubble-enriched space, $\Sigma_{k,h}^{\rm tn}$, was introduced in [21] for the Stokes equation with weakly imposed stress symmetry. Our newly developed tangential-normal continuous finite element space $\Sigma_{k,h}^{\rm tn}$ can also be used to discretize a distributional formulation of the Stokes equation, which will be explored elsewhere.
We use Nédélec elements $\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}\subset V_{0}^{{\rm curl\,}}$ for the tangential continuous vector space, and use the Riesz representation of the $L^{2}$-inner product to bring the abstract dual to a concrete function. Define $({\rm curl\,}\operatorname{div})_{h}:\Sigma_{h}^{\rm tn}\to\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}$ such that
(3) $(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})=\langle({\rm curl\,}\operatorname{div})_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h}\rangle,\quad\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}},$
@@
where $({\rm grad\,},\operatorname{mskw}\boldsymbol{x})\begin{pmatrix}\boldsymbol{v}\\ c\end{pmatrix}={\rm grad\,}\boldsymbol{v}+c\operatorname{mskw}\boldsymbol{x}$, and $\ell=k-1$ or $\ell=k$ is introduced to distinguish the first and second kind of Nédélec element. The lowest order, i.e. $k=1,\ell=0$, of the last three elements are illustrated in Fig. 1.
The finite element complex (4) is a discretization of the distributional ${\rm curl\,}\operatorname{div}$ complex:
@@
$H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})\xrightarrow{{\rm curl\,}\operatorname{div}}H^{-1}(\operatorname{div},\Omega)\xrightarrow{\operatorname{div}}H^{-1}(\Omega)\rightarrow 0,$
where $H^{-1}(\Omega):=(H_{0}^{1}(\Omega))^{\prime}$, and Sobolev spaces of negative order
@@
(5a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{k-1,h},\psi_{h}\in\mathring{\mathbb{V}}_{\ell+1,h}^{{\rm grad\,}},$
(5b) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{(k,\ell),h}^{{\rm curl\,}},$
where $b_{h}(\boldsymbol{\tau},\psi;\boldsymbol{v}):=(({\rm curl\,}\operatorname{div})_{h}\boldsymbol{\tau},\boldsymbol{v})+({\rm grad\,}\psi,\boldsymbol{v})$ and $({\rm curl\,}\operatorname{div})_{h}$ is a discretization of distributional ${\rm curl\,}\operatorname{div}$ operator, cf. (3).
We prove two discrete inf-sup conditions and thus obtain the well-posedness of (5a)-(5b) and optimal order convergence
@@
Furthermore, we apply hybridization techniques to (5a)-(5b) which is equivalent to a stabilization-free weak Galerkin method and to the $H({\rm grad\,}{\rm curl\,})$ nonconforming finite elements in [30, 48] for solving the quad curl problem. For other discretization of the quad-curl problem, we refer to the macro finite element method in [29], nonconforming finite element methods in [31, 42, 44], mixed finite element methods in [41, 11], decoupled finite element methods in [9, 47, 8], and references cited therein.
@@ 2. Distributional ${\rm curl\,}\operatorname{div}$ complex
In this section, we present the distributional ${\rm curl\,}\operatorname{div}$ complex and introduce the weak differential operator $({\rm curl\,}\operatorname{div})_{w}$ which can be defined on the tangential-normal continuous matrix functions.
@@ 2.1. Notation
Define the surface gradient operator as $\nabla_{F}:=\Pi_{F}\nabla$. For a scalar function $v$, define the surface ${\rm curl\,}$:
${\rm curl\,}_{F}v=\boldsymbol{n}\times\nabla v=\boldsymbol{n}\times\nabla_{F}v.$
For a vector function $\boldsymbol{v}$, the surface rot operator is defined as:
${\rm rot}_{F}\boldsymbol{v}:=(\boldsymbol{n}\times\nabla)\cdot\boldsymbol{v}=(\boldsymbol{n}\times\nabla_{F})\cdot\Pi_{F}\boldsymbol{v}=\boldsymbol{n}\cdot({\rm curl\,}\boldsymbol{v}),$
which represents the normal component of ${\rm curl\,}\boldsymbol{v}$.
Denote the space of all $3\times 3$ matrices by $\mathbb{M}$, and all trace-free/traceless $3\times 3$ matrices by $\mathbb{T}$. Define the deviation $\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\frac{1}{3}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I}$. Obviously for a scalar function $u$, $\operatorname{dev}(u\boldsymbol{I})=0$. For a vector $\boldsymbol{w}=(\omega_{1},\omega_{2},\omega_{3})^{\intercal}\in\mathbb{R}^{3}$, let
@@ 2.2. The ${\rm curl\,}\operatorname{div}$ complexes
Next we use the framework developed in [13] to present a Helmholtz decomposition of $H^{-1}({\rm curl\,}\operatorname{div},\Omega;\mathbb{T})$. Denote by
@@
Notice that only interior faces are included in the second term as $\boldsymbol{n}_{F}\cdot{\rm curl\,}\boldsymbol{v}=\operatorname*{rot}_{F}\boldsymbol{v}=0$ for $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega)$.
@@ 3.1. Finite element spaces
$\mathbb{B}_{k}^{\rm tn}(T;\mathbb{T})=\mathbb{P}_{k-1}(T)\otimes\mathrm{span}\{\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{i}\otimes\boldsymbol{t}_{i\ell}),\lambda_{\ell}\operatorname{dev}(\boldsymbol{n}_{j}\otimes\boldsymbol{t}_{j\ell}),\ell=0,\ldots,3\}.$
@@
The construction can be readily extended to arbitrary dimension $\mathbb{R}^{d}$ for $d\geq 2$. There are $d+1$ faces for a $d$-simplex. At each face, we have $d-1$ linearly independent traceless matrices $\{\boldsymbol{t}_{F,i}\otimes\boldsymbol{n}_{F}\}_{i=1}^{d-1}$ and in total $(d+1)(d-1)$ such matrices form a basis of $\mathbb{T}$. The basis of bubble functions is constructed vertices-wisely $\{\lambda_{\ell}\operatorname{dev}(\nabla\lambda_{i}\otimes\boldsymbol{t}_{i\ell})\}$ for $\ell=0,\ldots,d$ and $d-1$ different $i$ for each $\ell$. ∎
Remark 3.6.
Since the basis functions are expressed in terms of intrinsic geometric quantities—tangential and normal vectors along with barycentric coordinates—the defined finite element spaces are affine invariant, meaning the space remains unchanged under affine transformations. Thus, the traditional method of using a reference tetrahedron and transforming to the physical element can be applied, resulting in the same finite element space. In particular, curved elements can be handled using the Piola transform, as described in equations (3.76)-(3.77) of [36]. Extending the two-dimensional work of [4] to a distributional mixed finite element method specifically tailored for the quad-curl problem on domains with curved boundaries is an interesting topic for future research.
@@ 3.2. Distributional finite element complex
Notice that $\operatorname{div}_{h}$ is the $L^{2}$-adjoint of $-{\rm grad\,}$ restricted to $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. Again for a fixed triangulation $\mathcal{T}_{h}$, $\operatorname{div}_{h}$ is well defined which can be obtained by inverting the mass matrix of $\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}$. However $\{\operatorname{div}_{h}\}$ is not uniformly bounded as $h\to 0$ as $(\boldsymbol{v},{\rm grad\,}\psi)$ is in general not a bounded linear functional of $L^{2}$ unless $\boldsymbol{v}\in H(\operatorname{div},\Omega)$.